RXTE Helpdesk/FAQ RXTE What's New HEASARC Site Map


Feb 6th 1996.

We plan to attempt to turn the ASM on again on Feb. 20th. Full details on the status of the ASM and the turn-on strategy are provided by MIT .


Feb 2nd 1996.

The IOC period is now complete and we have transitioned to normal observations. Hercules X-1 was one of the final IOC targets, the observation was done at this time in order to observe at an appropriate phase of the 35 day period exhibited by this system.

Monitoring observations of the TOO target GRO J1744-28 are continuing and a short TOO observation of 4U1630-47 was also performed.

We will continue to provide information about XTE Science Operations on this WWW page from time to time.


Jan 24th 1996.

We are now approaching the end of the main checkout period and are preparing to concentrate on Guest Observer observations for AO1. Schedules for these observations will be appearing on our Web pages in the near future. A small number of calibration observations will be done during the first week of AO1 observations. Throughout the lifetime of the mission additional calibration observations will naturally also be done from time to time as required.

An additional Target of Opportunity observation of the Galactic Center transient source GRO J1744-28 is being planned for January 27th.


Jan 22th 1996.

Although we are still primarily carrying out calibration observations, the first XTE GO observations from AO1 are now being done. Regular monitoring of 3C 279 is currently being performed each day (proposal number 10354) for 500 seconds. The short length of each observation gives only a small impact on the calibration data.

Initial results from the XTE observation of GRO J1744-28 are reported in IAU Circular number 6291. Additional information on XTE Target of Opportunity observations is also available.


Jan 17th 1996.

Preparations are underway for tomorrow's observation of GRO J1744-28. Due to the brightness of this source we currently intend to point at an offset position in order to reduce the count rate to approximately twice that of the Crab. This will reduce dead time effects in the PCA caused by high count rates, especially during any bursts that may occur. In order to determine an appropriate pointing position the initial part of the observation will consist of a short series of scans across what is thought to be the actual source location.

After an initial change at launch, the PCA gain now appears to be very stable. This is consistent with our experience with this detector on the ground. Also, with its automatic gain control enabled, the HEXTE gain is stable to 0.25 %


Jan 16th 1996.

The first XTE observation specifically intended for science rather than calibration purposes was performed today when the PCA and HEXTE collected data on NGC 4945. As there is no pre-approved proposal to observe this source the observation is public domain and the data will be made available to the community on a similar timescale to other data collected during the IOC period.

The next science target to be observed will be GRO J1744-28 with observations scheduled to start late on January 18 and continue until early on January 19.

Detailed updates on the status of the ASM are being provided by MIT.


Jan 15th 1996.

The PCA and HEXTE experiments continue to perform well and we are accumulating additional calibration data.

It has been decided by the Project Scientist to perform a Target of Opportunity observation of the Seyfert 2 galaxy NGC 4945 during the IOC period (Jan. 16th). The decision to observe this source now was based on factors which included the facts that GRO (OSSE) and ASCA observations are currently being performed and that NGC 4945 is presently known to be bright.

The ASM, however, has experienced additional problems. The (azimuthal) ASM SSC2 experienced a high-count-rate turnoff on Friday night. It now appears that the failures in both SSC3 and SSC2 are due to internal high-voltage breakdown. One of the eight anodes was involved in SSC2 and two in SSC3. All other anodes appear to be unaffected. All three SSCs will be turned off in the interim while the data are being analyzed and possible recoveries are studied.

The ASM motor is still carrying out planned rotations and dwells and the rotation readouts are being used to checkout rotation algorithms. The sun-avoidance HV on-off commanding has been implemented and is in the command load for the current week.


Jan 11th 1996.

While the "Blizzard of 1996" has caused delays in updating this Web page XTE operations have continued in spite of the heavy snow. Members of the XTE team living close to GSFC managed to make their way into work but at least one person found themselves unexpectedly working a 36 hour shift until their replacement could arrive!

Targets that have now been observed by XTE also include SMC X-1, Cyg X-2 and MCG8-11-11. Real time software clearly showed the expected pulsations from SMC X-1 in the PCA data and off-line analysis indicates QPO in the Cyg X-2 data. Initial results from observations designed to measure instrument bore-sights show that the PCA and HEXTE are aligned to better than a few arc minutes.

Two of the SSCs appear to be working very well and the drive assembly is performing as expected. However, SSC 3 is having some problems which are being investigated. In addition, we are finding that pointing an SSC at the Sun causes a count rate which is high enough to switch off that SSC. We are therefore modifying our daily command loads so as to automatically turn off the high voltage supply to an SSC whenever the Sun will be in the field of view of the instrument.

Our initial model of the SAA has proved to be somewhat over-optimistic and we are now going to use a bigger region. Particle monitor data from HEXTE is being utilized in order to find the best contour to use for the SAA.

Upcoming observations include more pointings at MCG8-11-11, the Perseus cluster, and additional observations of the Crab to provide better information on dead time and instrument field of view.

Later in the month we also hope to be able to perform observations of the recently discovered Galactic center transient source GRO J1744-28.


Jan 6th 1996.

Today the PCA and HEXTE made their first observations of cosmic X-ray sources. First to be observed was the pulsar PSR 1509-58. Data were first received while the source was occulted by the Earth and, as the pulsar emerged from Earth occult, the increase in count rate as we detected the source could easily be seen on our real time display of the PCA light curve. Subsequent analysis clearly showed the presence of pulsations and HEXTE was able to measure the spectrum of the the pulsar out to beyond 60 keV.

Next to be observed was the Crab nebula. Initial analysis indicates that we are seeing count rates close to what were predicted before launch.

ASM scans were made and Sco X-1 was very obvious in real time data. The galactic center region was also observed.

Tonight and tomorrow very heavy snow falls are predicted for the Goddard area. However, we hope to be able to continue with satellite operations and proceed with further calibrations using the Crab nebula and our first observation of an AGN, MCG8-11-11.


Jan 5th 1996.

XTE has now made its first observations of X-ray sources! The high voltages to the ASM SSCs were turned on with the detectors rotating for several hours and a number of sources, including the Crab, were detected. Complete information on source positions and fluxes awaits calibration of detectors and verification of software.

The PCA and HEXTE detectors were also operated while viewing "empty" positions on the sky and these experiments continue to perform with no apparent problems.

We have now been given the go-ahead by the engineers to move to the next stage of the checkout phase. The spacecraft will operate with both high gain antennas (instead of one HGA and one omni antenna) and the Virtual Recorder has been repartitioned so as to increase the ratio of science to engineering data that is stored.

Extensive calibration observations can now proceed. The observations planned for Jan. 6 are:
PSR 1509-58 2 spacecraft orbits to verify accuracy of the spacecraft clock
Crab nebula 4 orbits to investigate instrument dead time and energy response
Crab nebula 3 orbits offset from the Crab by 1 degree to characterize instrument point spread function
Background 6 orbits with 5 separate pointing directions

The ASM will operate again and regions of the sky that are observed are expected to include the Galactic center.


Jan 4th 1996.

An extensive series of test pulse generator (TPG) experiments was run on the PCA with the EDS operating in both Standard Modes and using the Good Xenon modes. Tests to characterize the pulse height and pulse shape characteristics of HEXTE were also carried out. In preparation for turning on the high voltages to the ASM additional monitoring of SSC rotation was performed.

Tomorrow there will be an engineering meeting where it will be decided whether the XTE spacecraft is ready to enter the next phase of check out where the science instruments have priority.


Jan 3rd 1996.

All five proportional counter units of the PCA have now had high voltages turned on simultaneously and run for several hours. The detectors appear to be operating as predicted. The high voltage supplies to both HEXTE clusters were turned on last night and are now expected to remain on for the rest of the lifetime of XTE. Rotation tests of the ASM continue and no problems have been found.


Jan 2nd 1996.

The emphasis on operating XTE has, so far, been on checking out the spacecraft systems rather than the science instruments. As the engineers complete their tests more time will be made available for checking that the science instruments are working correctly and then progressing to calibration observations.

As of today, high voltage has been turned on at normal operating levels for three out of the five PCA proportional counter units and both HEXTE clusters. Both instruments appear to be healthy and showing close to expected behavior. HEXTE rocking tests have also been successful. The ASM high voltage will be turned on later but the drive has been turned on and the SSCs have been seen to rotate following the commanded plan.


Please send questions/problem reports to: xtehelp@athena.gsfc.nasa.gov
HEASARC/Office of Guest Investigator Programs


For questions about scheduling of RXTE observations please send email to xteplan@athena.gsfc.nasa.gov If you have other questions about RXTE, please send email to one of our help desks.

This page is maintained by the RXTE SOF and was last modified on Thursday, 03-Jun-2004 10:36:08 EDT.

Responsible NASA Official: Phil Newman
Web Curator: Robin Corbet

Privacy Policy and Important Notices.
Last Revised: Thursday, 03-Jun-2004 10:36:08 EDT